Your computer manufacturing firm must purchase 12,000 keyboards from a supplier. One supplier demands a payment of  $144,000 today plus $12 per keyboard payable in one year. Another supplier will charge $25 per keyboard, also payable in one year. The risk-free interest rate is 9%. 

a. What is the difference in their offers in terms of dollars today?  Which offer should your firm take? 

b. Suppose your firm does not want to spend cash today. How can it take the first offer and not spend $144,000 of its own cash today?
Answer #1

A)

step 1) pv= 144,000+ 12*12,000/ (1+ 0.09)= 276,110.09174312 or 276,110

step 2) pv= 25*12,000/ (1+0.09)= 275,229.35779817 or 275,229

step 3) 276,110- 275,229= 881 answer


The firm should take second supplier's offer.


B) It can borrow the required payment today at 9%.

answered by: Andrew San Andres